OpenStudy (imtant):

How to determine the value of young's modulus of elasticity if a steel bar, 40mm*10mm section and 2 m length is subjected to an axial compressive load of 40 kN. The contraction of the length is found to be 1 mm and the increase in 40 mm width is 0.006m?

OpenStudy (imtant):

Given the formula is E= PL/ A delta.E for young's modulus of elasticity. P for 40 kN.

OpenStudy (imtant):

@Jack1

OpenStudy (imtant):

Okay, but may I know why delta L is in negative?

OpenStudy (imtant):

And for the A (cross section area) I got 0.0004 m^2.

OpenStudy (imtant):

So for the answer I got 200 Gpa. Is it same with you?

OpenStudy (jack1):

so F = compressive load of 40,000N length has reduces due to compression by 1 mm width has increased by 6mm using your equation E= PL/ A delta.E P = 40000 N L = 2m ? delta e = extension? A = ... initial or new now we have the different width... ?

OpenStudy (jack1):

so ur equation was 40000 x 2 / 0.04 x 0.01 x ...0.001?

OpenStudy (jack1):

what did u use as delta e?

OpenStudy (imtant):

Yes , it is. For delta or E?

OpenStudy (jack1):

E= PL/ A delta. so A delta = delta A ? = change in area?

OpenStudy (imtant):

\[E= \frac{ PL }{ A \delta } \]

OpenStudy (imtant):

I'm just using 0.0004m.

OpenStudy (jack1):

0.04 x 0.01 = 0.0004 m^2 initial new area = 0.039 x 0.046 = 0.001794 m^2

OpenStudy (jack1):

so delta area = 0.001394m^2 ... yeah?

OpenStudy (jack1):

sorry dude, i think i'm just muddying the waters, it's been too long since i learned physics i think

OpenStudy (jack1):

@dan and @ganeshie8 do you guys have a moment please?

OpenStudy (imtant):

It's okay Jack.

OpenStudy (imtant):

I think I will just using 0.001 for delta. And area are 0.0004 from (0.040*0.010).

OpenStudy (imtant):

Anyways Thanks Jack. :)

OpenStudy (imtant):

\[\delta = elongation.\]